Danny drove halfway from town P to town Q at an average speed of 60km/h. He wants to average 80km/h for the entire journey. What will Danny's average speed for the second part of his journey?
Answer by Alan3354(69443)

Danny drove halfway from town P to town Q at an average speed of 60km/h. He wants to average 80km/h for the entire journey. What will Danny's average speed for the second part of his journey?
---------------
Avg speed for a round trip (or of equal distancts) is
2*r1*r2/(r1+r2) (r1 & r2 are the 2 speeds)
(I can show the derivation if needed)
2*60*r2/(60+r2) = 80
60*r2 = 40*60+r2) = 40*r2 + 2400
20r2 = 2400
r2 = 120 km/hr